OpenStudy (anonymous):

If the point (a,3) lies on the graph of the equation 5x + y = 8, then a=

OpenStudy (joannablackwelder):

Since "a" is in the position of x in your coordinate pair, you can plug in 3 for y. The equation is then 5x + 3 = 8. Solving for x gives 5x=5, so x=1. As I said earlier, a would be the value of x when y is 3, so a=1.

OpenStudy (anonymous):

thanks joanna 1 more Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (joannablackwelder):

Sure. Start out with the point slope equation. y-y1=m(x-x1) Plug in 3/5 for m, 2 for x1, and -4 for y. This gives y-(-4)=3/5(x-2) Reduce it to y+4=3/5x-6/5 And solve for y y=3/5x-6/5-20/5 y=3/5x-26/5 Hope that helps!

OpenStudy (anonymous):

that did !! thanks ! The equation of the line through point (- 4, 0) and parallel to the y-axis is??

OpenStudy (joannablackwelder):

A line that is paralell to the y axis is a line that is x=a constant. Since we know that one possible value of x is -4, the line must be x=-4.

OpenStudy (anonymous):

ok perfect what about this one? this stuff is so tough ! The slope of the line containing the points (-5, 3) and (-2, 1) is

OpenStudy (joannablackwelder):

Slope is (y2-y1)/(x2-x1). You can pick either point to be point 1 or point 2. Try seeing if that helps and let me know what you get slope to be.

OpenStudy (anonymous):

i got -.666 repeating

OpenStudy (joannablackwelder):

Perfect! Good job. You can leave it as a fraction of -2/3 if you want to. :)

OpenStudy (anonymous):

ok can you stay on please jic i need more help?

OpenStudy (anonymous):

Given a dilation with the origin O (0, 0), by observation determine the scale factor "K." DO, K = (5, 0) (15, 0) What is the scale factor K? like now haha

OpenStudy (joannablackwelder):

I'll try to be on as much as I can. If you would like a formal lesson, I am a professional tutor and can do online tutoring with video chat and a whiteboard at $30/hr. Let me know if you are interested. :)

OpenStudy (joannablackwelder):

A dilation is a factor that changes x and y by the same constant. So, since (5,0) is your original point and your dilated point is (15,0), that can be written as (3*5,3*0). Thus the K=3.

OpenStudy (anonymous):

ok what about this is it true or false? 0, -180 is the same rotation as 0, 180

OpenStudy (joannablackwelder):

0 degree rotation puts you on the x axis facing right (assuming normal conventions). -180 degree rotation puts you on the x axis but facing left. These are not the same. It is false.

OpenStudy (anonymous):

thats what i thought, because its negative not positive right?

OpenStudy (joannablackwelder):

No, positive 180 degrees and 0 degrees are not the same either. 360 degrees and 0 degrees or -360 degrees and 0 degrees are the same though.

OpenStudy (anonymous):

T : (x, y) (x + 4, y + 3) T-1: (x, y) (x - 4, y - 3) (-4x, -3y) (x + 3, y + 2) ok how about this. i chose the last one

OpenStudy (joannablackwelder):

I don't understand the question.

OpenStudy (anonymous):

For the transformation T, what is T-1?

OpenStudy (joannablackwelder):

I think what it means is that for the translation T: (x,y)->(x+4,y+3). So for T-1, that translation should be one less than the T translation., putting it at (x+3,y+2). Disclaimer: I'm still not sure if I am understanding the problem here. :/

OpenStudy (anonymous):

yup we got it

OpenStudy (joannablackwelder):

:)